Jabu had a few marbles. Today he played and doubled his number of marbles. Then Thabo gave him three marbles for free. Jabu has 21 marbles now. How many did he have before he started playing? Solve this problem using a suitable model.
1
Expert's answer
2020-07-07T16:42:42-0400

Solution:

Jabu had a few marbles.


"x."

Today he played and doubled his number of marbles.


"2x."

Thabo gave him three marbles for free.


"2x+3."

Jabu has 21 marbles now.


"2x+3=21."

To find out how many marbles there were initially, you need to solve this equation.


"2x+3=21;""2x=21-3=18;""x=18\/2=9."

Answer:

Jabu had 9 marbles before the game.